# Genomic Language Models: Opportunities and Challenges

### v1

@@ Abstract
Large language models (LLMs) are having transformative impacts across a wide range of scientific fields, particularly in the biomedical sciences. Just as the goal of Natural Language Processing is to understand sequences of words, a major objective in biology is to understand biological sequences. Genomic Language Models (gLMs), which are LLMs trained on DNA sequences, have the potential to significantly advance our understanding of genomes and how DNA elements at various scales interact to give rise to complex functions. In this review, we showcase this potential by highlighting key applications of gLMs, including fitness prediction, sequence design, and transfer learning. Despite notable recent progress, however, developing effective and efficient gLMs presents numerous challenges, especially for species with large, complex genomes. We discuss major considerations for developing and evaluating gLMs.
@@ INTRODUCTION
[style=box] At a high level, a language model is trained to learn the conditional probability distribution of the form $\mathbb{P}[X_{i}\mid X_{-\text{Masked}}]$ for $i\in\text{Masked}$ (in Masked Language Modeling, MLM) or $\mathbb{P}[X_{k}\mid X_{1:k-1}]$ (in Causal Language Modeling, CLM), where $X=(X_{1},X_{2},\ldots)$ denotes a sequence of “tokens” (e.g., nucleotides or amino acids) and “Masked” denotes a collection of masked positions. The key to recent advances in NLP is that, instead of fitting a parametric distribution that one designs by hand, one lets the dataset speak for itself and fit more complex models as more data are observed. This nonparametric density estimation is achieved by leveraging deep learning. Figure 1 depicts the masked language modeling framework for DNA. While the model is trained to predict the nucleotide at each masked site using information from unmasked sites, it will learn position-specific contextual representation (called embedding, a high-dimensional vector in $\mathbb{R}^{n}$), which then gets converted into a probability distribution over {A,C,G,T}. These embeddings and probability distributions, both of which are position-specific, can be applied to many problems in genomics.
@@
Figure 1: Training and applications of gLMs. The schematic on the left-hand side illustrates MLM training. The log-likelihood ratio (LLR) between two alleles (specifically, $\log[\mathbb{P}(X_{i}=a\mid X_{-i})/\mathbb{P}(X_{i}=b\mid X_{-i})]$) is a good unsupervised predictor of fitness or deleteriousness (Fitness prediction). New sequences can be generated by sampling from the learned probability distribution (Sequence design). A vector representation, called embedding, of each token in the input sequence can be extracted and adapted for different downstream tasks (Transfer learning).
@@
Figure 2: Application examples. (a) gLM GPN’s predicted logo plot (top) at a promoter, highlighting a motif (bottom logo) that matches a putative functional TFBS. (b) Correlation between minor allele frequency (MAF) and GPN score (log-likelihood ratio). (c) A gLM can be prompted with different control tags to design promoter sequences driving high or low expression in a given cell type. (d) Visualization of GPN embeddings for different classes of genomic windows, illustrating that the learned representations contain useful information such as gene regions.
@@
The general language model framework is summarized in Box 1. Below, we elaborate on three main application areas of gLMs: Fitness prediction, Sequence design, and Transfer learning.
Fitness prediction
@@
An intriguing application of gLMs is the unsupervised prediction of the fitness (specifically, deleteriousness) of genetic variants. The underlying idea is that reference genomes, typically derived from healthy individuals, are relatively depleted of deleterious variants. Consequently, models trained on these data are predisposed to assigning lower probabilities to harmful variants. This observation underpins the strategy of using the log-likelihood ratio (LLR) between two alleles (e.g., $\log[\mathbb{P}(X_{i}=a\mid X_{-i})/\mathbb{P}(X_{i}=b\mid X_{-i})]$ in the MLM setting described in Box 1) to estimate their relative fitness. A significant benefit of this approach is its independence from supervised labels, such as whether a variant is disease-causing, which are often limited and subject to a wide range of biases.
@@
Fitness prediction using the LLR was initially introduced in the context of protein sequence models, leading to outstanding results in predicting the effects of missense variants [10, 11, 6, 12]. Expanding this approach, genome-wide fitness prediction using a gLM was first undertaken by GPN [13], achieving state-of-the-art results in the model plant Arabidopsis thaliana. To illustrate how a gLM might be able to predict fitness, we note that gLMs can learn transcription factor binding site (TFBS) motifs, understanding which positions are under constraint and which are not (Figure 2a). In addition, GPN’s LLR score is correlated with allele frequencies in natural Arabidopsis thaliana populations, even though the model was only trained on a single genome from this species (Figure 2b). Subsequently, AgroNT [14] and PlantCaduceus [15] have also obtained excellent results in other plant species. For the human genome, however, the LLR from the Nucleotide Transformer (NT) [16] fell short of existing baselines. Meanwhile, GPN-MSA [17], leveraging a whole-genome multiple sequence alignment (MSA) across diverse vertebrate species, was able to attain state-of-the-art performance (see Learning objective for further MSA considerations). It should be noted that the observed nucleotide distribution is driven not only by fitness but also by mutational biases; explicitly incorporating this information into fitness prediction is a promising avenue of future research.
@@
There are two main kinds of variant effect predictors in genomics: fitness predictors, including gLMs but also conservation scores, and activity predictors, such as the gene expression predictor Enformer [18] or the splicing predictor SpliceAI [19]. These two kinds of models are related in the sense that if a variant changes fitness, it does so via a change in activity in some context (e.g., change in transcription of a certain gene during limb development), ultimately affecting a high-level trait (e.g., polydactyly). Fitness models cover all possible mechanisms and contexts that affect the overall organismal fitness, while activity models reflect only those they are explicitly trained on (some data, such as protein expression in the developing human brain, are just hard to obtain). On the other hand, activity models can nominate a specific mechanism and context through which a variant acts, while fitness models are more of a black box.
@@
With regards to functional variant prioritization, there are some additional considerations. An activity model would not be able to prioritize between two variants that induce a similar gene expression fold change but in two different genes, which might have completely different impacts on high-level traits. On the other hand, a trait not under detectable selection could still be of scientific or medical interest. In this case, a fitness model would have limited power to prioritize variants affecting it, especially if they have small effect sizes, as is the case in complex trait GWAS. However, while a gLM’s LLR might not have high power in this setting, gLM’s learned embeddings (Box 1) could still provide value with additional supervision on labeled data [20].
@@
Designing novel biological sequences is of great interest to both the academic and industry research communities due to its immense potential in drug discovery and delivery; agricultural improvement; bioremediation; and the development of biological research tools. We here describe sequence generation with a CLM (Box 1) as it is the most common approach (see Learning objective for generation with MLMs). Specifically, the sequence generation task is decomposed into a series of next-token prediction problems. Starting with a given sequence fragment (referred to as prompts [21], or control tags [22]), the language model can predict the next token recursively and generate a whole new sequence. pLMs have been shown to be powerful tools for protein design [22, 23, 24, 25]. Going beyond coding sequences, designing non-coding sequences is also crucial due to its applications such as gene and cell therapies [26], as well as synthetic biology [27]. Recent work has explored the use of gLMs for de novo DNA sequence generation.
@@
Additionally, gLMs can be potentially used to design organized, functional DNA sequences at the chromosome or genome-scale. Recently, two gLMs, MegaDNA and EVO, have explored such design tasks for prokaryote genomes [29, 27]. EVO pretrained model was used to generate 20 sequences of size about 650 Mbp. The generated sequences were found to have realistic coding sequence density, protein sequences with predicted secondary structure and globular folds, as well as plausible tRNA sequences. MegaDNA was used to generate full bacteriophage genomes up to 96 kbp. Apart from validating coding sequences, the author further identified functional regulatory elements including promoters and ribosome binding sites in the generated sequences. Yet, such mega-scale DNA sequence design tasks remain challenging. The generated sequences by EVO were found to lack highly conserved marker genes that typically exist in functional prokaryote genomes, and the predicted protein structures have limited matches to natural protein databases. A recent independent evaluation [30] revealed that the sequence composition of MegaDNA-generated genomes is still largely dissimilar to natural genomes. Therefore, further work is needed to refine the methods to enable de novo design of fully functional genomes with gLM.
@@
Neural networks trained to predict annotations from functional genomics experiments have been widely utilized to interpret the functions of genomic elements. A significant application has been predicting variant effects on molecular phenotypes, such as gene expression [31, 32, 33, 34, 18] and splicing [19, 35]. The ability of neural networks to interpret complex interactions between genomic sites has made them essential tools for tackling these important problems, but suitable training data are often difficult to collect and consequently limited. To generalize on prediction tasks, models need to develop an understanding of genomic grammar (such as sequence motifs and epistatic interactions), which requires substantial data and computation. To overcome the limitations of insufficient data for individual tasks, developers have employed transfer learning methods — techniques that leverage knowledge gained from training models on one task to improve performance on related tasks. Specifically, most neural networks trained to predict functional annotations have been trained to predict a wide array of annotations simultaneously, forcing these models to learn a single unifying representation. This, in turn, has improved their generalization performance.
@@
[style=box] For NLP models to generalize on most tasks (including typical tasks, such as sentiment analysis, question answering, and part-of-speech tagging, to name only a few), models need to understand grammar and meaning. However, data specific to these tasks are limited. Utilizing LLMs trained on raw text data (sourced from articles, books, and websites) for transfer learning has enabled breakthrough progress on these problems [36]. Today, virtually every state-of-the-art NLP model is adapted from a LLM.
@@
Language models may also be utilized for transfer learning. See Box 2 for the general concept of transfer learning in NLP. One technique is feature extraction: while learning to predict the context-dependent distribution of nucleotides, gLMs transform input genomic sequences into intermediate vector representations (Box 1). These representations may distill relevant information, and, therefore, be utilized as features for another model. For example, visualization of GPN embeddings reveals that, without any supervision, the model has learned to distinguish different classes of genomic elements such as coding sequence and untranslated regions [13] (Figure 2d). Another way to utilize language models for transfer learning is to use them as pretrained models: that is, to continue training them on downstream tasks. This technique is called fine-tuning. Fine-tuned models develop representations that synthesize knowledge from both tasks, which may improve their generalization performance on the downstream task.
For example, SegmentNT [38], a fine-tuned version of a gLM called Nucleotide Transformer (NT) [16], was recently developed to annotate gene regions and cis-regulatory elements in the human genome at base-pair resolution. Without the MLM pretraining step, the obtained performance was substantially worse. Furthermore, the authors evaluated the accuracy of the model on species not seen during training and how it correlates with phylogenetic distance. AgroNT [14], another model of the NT family, was pretrained on diverse plant species and then fine-tuned to predict chromatin accessibility and gene expression on select crop species.
Two recent studies evaluated several gLMs in prediction tasks in the human genome and found that they did generally not outperform non-gLM baselines [39, 40]. These results were based on frozen embeddings; evaluating full fine-tuning would provide additional insights. While gLMs are already well suited to demonstrate the value of transfer learning in less-studied organisms, further innovation may be required for them to offer significant value in human genetics, where high-quality labeled data and carefully crafted models already exist. An important question is how far the scaling hypothesis holds for gLMs, i.e., how much increasing unlabeled data and computation will keep improving model performance.
@@
LookingGlass [46] Bacteria + archaea CLM RNN nucleotide-level Metagenomic sequences from diverse environments rather than assembled genomes are used for training.
@@
SpliceBERT [52] Human + 71 veterbrate genomes MLM Transformer nucleotide-level Only RNA Transcripts are used in training.
species LM [53] 1500 fungal genomes MLM Transformer overlapping k-mer Only 5’ and 3’ UTR regions are used in training: the 5’ species LM and 3’ species LM.
@@
UTR-LM [58] Human + 4 vertebrate genomes MLM Transformer nucleotide-level Only 5’ UTR regions are used in training. Trained also to predict mRNA minimum free energy and secondary structures calculated by ViennaRNA [59].
@@
AgroNT [14] 48 plant genomes focusing on edible plant species MLM Transformer BPE
@@
ChatNT [63] DNA sequences + English intructions CLM Transformer overlapping k-mer Combines the pretrained gLM NT [16] and the English LM Vicuna [64].
@@
CD-GPT [65] Human + multi-species CLM Transformer BPE DNA/RNA/Protein multi-modal pretraining.
Table 1: A summary of existing gLMs. An overview of various gLMs is provided, highlighting their pretraining datasets, tasks, architectures, tokenization methods, and unique features. The models are listed in the order of their public release dates.
@@
The performance of a machine learning model is significantly influenced by both its architecture and its training data. Various model architectures such as convolutional neural networks (CNNs), Transformers, and state-space models have been successfully adapted to a wide range of domains, including natural language, images, audio, proteins, and genomics. However, selecting suitable training data requires a deep understanding of the specific domain, especially in genomics where there is no universally accepted, curated dataset comparable to those in NLP (e.g., the Pile [66, 67]) or proteins (e.g., UniProt [4]).
@@
Another key question is how to ensure that the amount of data is enough. It is likely that a single genome might not be enough to train a large model, especially if non-functional regions are downsampled or downweighted. One approach is to add sequence variants from the same species [16]. However, in many cases, e.g., non-African human populations, there is relatively little variation between individuals. A more common approach is to train across multiple species (Table 1), as typically done for pLMs. As species become more distant, the grammar of the non-coding genome starts to change, faster so than the grammar of protein-coding regions. One proposed approach is to explicitly add a species identifier as an extra input to the model [53]. Notwithstanding, it is plausible that a large enough model, with enough genomic context, might be able to naturally model distant genomes, similarly to how LLMs handle multilingual datasets.
As mentioned earlier, in prokaryotes, there exist models (MegaDNA and EVO) that take an entire genome as context [29, 27]. This is currently infeasible for eukaryotes, and therefore leads to the question of how to partition the genome into context windows to be separately modeled. Many interactions are restricted to nearby positions, such as transcription factor binding site motifs, motivating the development of models with a relatively small context ($<$ 6 kb) (Table 1). However, there are obvious long-range interactions, such as between exons of the same gene or between enhancers and promoters (up to 1 Mb) [70]. Such long context lengths introduce computational and statistical challenges, and efforts have been made to overcome them [28, 29, 27, 20]. Regardless of the chosen context length, it is still not easy to partition the genome into independent units (similarly to how proteomes are separated by protein). For instance, the enhancer of a gene can be located inside the intron of another gene [70].
@@
Despite their strengths, Transformer models face several challenges unique to genomic modeling. One significant issue is that Transformers have weak or no inductive biases regarding the locality of interactions [75, 41], making them less data-efficient at modeling local motifs such as TFBS. While this can be partially mitigated with CNN-Transformer hybrid architectures such as LOGO [47], further research is needed to enhance this aspect.
@@
To overcome the quadratic scaling issues of self-attention, various state-space models (SSMs) [78, 79, 80] have been proposed for gLMs as efficient alternatives to Transformers, offering nearly linear scaling with sequence length. HyenaDNA [28], based on the Hyena Hierarchy [79], can support input contexts as long as 1 million nucleotides. EVO [27], a hybrid model combining Hyena and Transformer architectures, is pretrained with 8 kb sequences and later fine-tuned with 131 kb sequences during the context extension stage. Caduceus [20], built on the Mamba-based SSM [80], is trained on 131 kb sequences while incorporating reverse-complementarity equivariance.
@@
To reduce input sequence length and model longer context, both $k$-mer and byte-pair encoding [85] tokenizations create artificially defined nucleotide vocabularies larger than the natural nucleotide vocabularies of {A, C, G, T}. On the other hand, single-nucleotide tokenization simplifies model interpretation and attribution, and enhances the model’s ability to handle genomic variations more effectively.
Several modifications to the training objective have been explored to provide additional signal and boost performance. For instance, GPN-MSA [17] enhances MLM training on the human reference genome with a whole-genome MSA [86, 87] of vertebrate species, leveraging conservation across related species for additional context. A limitation is that whole-genome MSAs have only been generated for certain species, and might require further development to be effective in plants [88]. Similarly, Species LM [53] directly integrates species information by assigning a dedicated token for each yeast species and appending the species token to the input sequence during training and inference. Training on nucleotide sequences has been expanded to enable cross-talk with additional modalities such as epigenetics [43, 44], RNA [65], proteins [65], and natural language [63].

### v2

@@ Abstract
Large language models (LLMs) are having transformative impacts across a wide range of scientific fields, particularly in the biomedical sciences. Just as the goal of Natural Language Processing is to understand sequences of words, a major objective in biology is to understand biological sequences. Genomic Language Models (gLMs), which are LLMs trained on DNA sequences, have the potential to significantly advance our understanding of genomes and how DNA elements at various scales interact to give rise to complex functions. To showcase this potential, we highlight key applications of gLMs, including functional constraint prediction, sequence design, and transfer learning. Despite notable recent progress, however, developing effective and efficient gLMs presents numerous challenges, especially for species with large, complex genomes. Here, we discuss major considerations for developing and evaluating gLMs.
@@ INTRODUCTION
[style=box] At a high level, a language model is trained to learn the conditional probability distribution of the form $\mathbb{P}[X_{i}\mid X_{-\text{Masked}}]$ for $i\in\text{Masked}$ (in Masked Language Modeling, MLM) or $\mathbb{P}[X_{k}\mid X_{1:k-1}]$ (in Causal Language Modeling, CLM), where $X=(X_{1},X_{2},\ldots)$ denotes a sequence of “tokens” (e.g., nucleotides or amino acids) and “Masked” denotes a collection of masked positions. The key to recent advances in NLP is that, instead of fitting a simple parametric model of context dependency that one designs by hand, one lets the data speak for themselves and fit more complex models as more data are observed, by leveraging powerful deep learning architectures. Figure 1 depicts the language modeling framework for DNA. While the model is trained to predict the nucleotide at each masked site using information from unmasked sites, it will learn position-specific contextual representation (called embedding, a high-dimensional vector in $\mathbb{R}^{n}$), which then gets converted into a probability distribution over {A,C,G,T}. These embeddings and probability distributions, both of which are position-specific, can be applied to many problems in genomics.
@@
Figure 1: Training and applications of gLMs. The schematic on the left-hand side illustrates gLM training. The log-likelihood ratio (LLR) between two alleles (specifically,$\log[\mathbb{P}(X_{i}=a\mid X_{-i})/\mathbb{P}(X_{i}=b\mid X_{-i})]$) is a good unsupervised predictor of functional constraint (Functional constraint prediction). New sequences can be generated by sampling from the learned probability distribution (Sequence design). A vector representation, called embedding, of each token in the input sequence can be extracted and adapted for different downstream tasks (Transfer learning).
@@
Figure 2: Application examples. (a) gLM predicted logo plot (top) at a promoter, highlighting a motif (bottom logo) that matches a putative functional TFBS. (b) Correlation between variant minor allele frequency (MAF) and gLM score (log-likelihood ratio). (c) A gLM can be prompted with different control tags to design promoter sequences driving high or low expression in a given cell type. (d) Visualization of gLM embeddings for different classes of genomic windows, illustrating that the learned representations contain useful information such as gene regions. Note: Panels a,b,d were generated using the GPN model.
@@
The general language model framework is summarized in Box 1. Below, we elaborate on three main application areas of gLMs: Functional constraint prediction, Sequence design, and Transfer learning.
Functional constraint prediction
@@
An intriguing application of gLMs is the prediction of functional constraint on a genomic locus without any supervision on the task. A significant benefit of this approach is its independence from labels, such as whether a variant is disease-causing, which are often limited and subject to biases. The underlying idea is that reference genomes, typically derived from healthy individuals, are relatively depleted of deleterious variants. Consequently, models trained on these data are predisposed to assigning lower probabilities to harmful variants. This observation underpins the strategy of using the log-likelihood ratio (LLR) between two alleles (i.e., $\log[\mathbb{P}(X_{i}=a\mid X_{-i})/\mathbb{P}(X_{i}=b\mid X_{-i})])$) to estimate their relative fitness.
@@
Functional constraint prediction using the LLR was initially introduced in the context of protein sequence models, leading to outstanding results in predicting the effects of missense variants [10, 11, 6, 12]. Expanding this approach, genome-wide functional constraint prediction using a gLM was first undertaken by GPN [13], achieving state-of-the-art results in the model plant Arabidopsis thaliana. To illustrate how a gLM might be able to predict functional constraint, we note that gLMs can learn transcription factor binding site (TFBS) motifs, understanding which positions are under constraint and which are not (Figure 2a). In addition, GPN’s LLR score is correlated with allele frequencies in natural Arabidopsis thaliana populations, even though the model was only trained on a single genome from this species (Figure 2b). Subsequently, AgroNT [14] and PlantCaduceus [15] have also obtained excellent results in other plant species. For the human genome, however, the LLR from the Nucleotide Transformer (NT) [16] fell short of existing baselines. Meanwhile, GPN-MSA [17], leveraging a whole-genome multiple sequence alignment (MSA) across diverse vertebrate species, was able to attain state-of-the-art performance (see Learning objective for further MSA considerations). It should be noted that the observed nucleotide distribution is driven not only by functional constraint but also by mutational biases; explicitly incorporating this information into functional constraint prediction is a promising avenue of future research.
For a single nucleotide polymorphism (SNP), the LLR can be computed in a single query to an MLM with the variant position masked, but in two queries to a CLM on the reference and alternate sequences. A CLM can just as easily handle multiple substitutions, insertions and deletions, while an MLM must resort to a more expensive pseudo-LLR [18, 12]. Scores other than the LLR have been proposed for functional constraint prediction, such as the distance in embedding space [16, 14] or the change in nucleotide probabilities in positions around a mutation [19]. While the LLR is widely used in both the pLM and gLM communities, it is important to understand better in which contexts these alternative scores are useful.
@@
There are two main kinds of variant effect predictors in genomics: functional constraint predictors, including gLMs and traditional conservation scores [20, 21], and activity predictors, such as the gene expression predictor Enformer [22] or the splicing predictor SpliceAI [23]. These two kinds of models are related in the sense that if a variant at a locus is under selection, it induces a change in activity in some context (e.g., change in transcription of a certain gene during limb development), ultimately affecting a high-level trait (e.g., polydactyly). Functional constraint models cover all possible mechanisms and contexts that affect the overall organismal fitness, while activity models reflect only those they are explicitly trained on (some data, such as protein expression in the developing human brain, are just hard to obtain). On the other hand, activity models can nominate a specific mechanism and context through which a variant acts, while functional constraint models do not offer a mechanistic interpretation.
@@
With regards to functional variant prioritization, there are some additional considerations. An activity model typically gives similar scores to two variants that induce a similar expression fold-change but in different genes, even if there is a vast difference in physiological tolerance to their expression levels. On the other hand, a trait not under detectable selection could still be of scientific or medical interest. In this case, a functional constraint model would have limited power to prioritize variants affecting it, especially if they have small effect sizes, as is the case in complex trait GWAS. However, while a gLM’s LLR might not have high power in this setting, gLM’s learned embeddings (Box 1) could still provide value with additional supervision on labeled data [24].
@@
Designing novel biological sequences is of great interest to both the academic and industry research communities due to its immense potential in drug discovery and delivery; agricultural improvement; bioremediation; and the development of biological research tools. We here describe sequence generation with a CLM (Box 1) as it is the most common approach (see Learning objective for generation with MLMs). Specifically, the sequence generation task is decomposed into a series of next-token prediction problems. Starting with a given sequence fragment (referred to as prompts [25], or control tags [26]), the language model can predict the next token recursively and generate a whole new sequence. pLMs have been shown to be powerful tools for protein design [26, 27, 28, 29]. Going beyond coding sequences, designing non-coding sequences is also crucial due to its applications such as gene and cell therapies [30], as well as synthetic biology [31]. Such design tasks have previously been addressed using supervised activity models [32, 33, 34], but more recently several works have explored the use of gLMs to tackle this challenge as described below.
@@
Additionally, gLMs can be potentially used to design organized, functional DNA sequences at the chromosome or genome-scale. Recently, two gLMs, MegaDNA and EVO, have explored such design tasks for prokaryote genomes [36, 31]. EVO was used to generate 20 sequences of size about 650 Mbp. The generated sequences were found to have realistic coding sequence density, protein sequences with predicted secondary structure and globular folds, as well as plausible tRNA sequences. MegaDNA was used to generate full bacteriophage genomes up to 96 kbp. Apart from validating coding sequences, the author further identified functional regulatory elements including promoters and ribosome binding sites in the generated sequences. Yet, such mega-scale DNA sequence design tasks remain challenging. The generated sequences by EVO were found to lack highly conserved marker genes that typically exist in functional prokaryote genomes, and the predicted protein structures have limited matches to natural protein databases. A recent independent evaluation [37] revealed that the sequence composition of MegaDNA-generated genomes is still largely dissimilar to natural genomes. Therefore, further work is needed to refine the methods to enable de novo design of fully functional genomes with gLMs.
@@
Neural networks trained to predict annotations from functional genomics experiments have been widely utilized to interpret the functions of genomic elements. A significant application has been predicting variant effects on molecular phenotypes, such as gene expression [38, 39, 40, 41, 22] and splicing [23, 42]. The ability of neural networks to interpret complex interactions between genomic sites has made them essential tools for tackling these important problems, but suitable training data are often difficult to collect and consequently limited. To generalize on prediction tasks, models need to be capable of identifying the broad set of functionally important sequence elements, which may require substantial data and computation. To overcome the limitations of insufficient data for individual tasks, developers have employed transfer learning methods — techniques that leverage knowledge gained from training models on one task to improve performance on related tasks. Specifically, most neural networks trained to predict functional annotations have been trained to predict a wide array of annotations simultaneously, forcing these models to learn a single unifying representation. This, in turn, has improved their generalization performance.
@@
[style=box] For NLP models to generalize on most tasks (including typical tasks, such as sentiment analysis, question answering, and part-of-speech tagging, to name only a few), models need to understand grammar and meaning. However, data specific to these tasks are limited. Utilizing LLMs trained on raw text data (sourced from articles, books, and websites) for transfer learning has enabled breakthrough progress on these problems [43]. Today, virtually every state-of-the-art NLP model is adapted from an LLM.
@@
Language models may also be utilized for transfer learning. (See Box 2 for a discussion of the utility of transfer learning for NLP.) One technique is feature extraction: while learning to predict the context-dependent distribution of nucleotides, gLMs transform input genomic sequences into intermediate vector representations (Box 1). These representations may distill relevant information, and, therefore, be utilized as features for another model. For example, visualization of gLM embeddings reveals that, without any supervision, the model has learned to distinguish different classes of genomic elements such as coding sequence and untranslated regions [13] (Figure 2d). Embeddings from different layers can provide information useful for different tasks [45]. Another way to utilize language models for transfer learning is to use them as pretrained models: that is, to continue training them on downstream tasks. This technique is called fine-tuning. Fine-tuning a pretrained neural network on a task implicitly regularizes its parameters such that the network’s predictions synthesize knowledge from both tasks. As a result, pretraining neural networks tends to improve their generalization performance on downstream tasks. In recent work, SegmentNT, a model developed by fine-tuning the Nucleotide Transformer (NT) gLM [16] to the task of annotating genes and cis-regulatory elements, achieved state-of-the-art performance on this task [46]. Utilizing a pretrained model was shown to be critical to its success. Similarly, AgroNT [14], another model of the NT family, was pretrained on diverse plant species and then fine-tuned to predict chromatin accessibility and gene expression on select crop species. DNABERT-S [47] applied contrastive learning with pretrained DNABERT-2 [48] embeddings to perform metagenomics binning. IsoFormer [49] is an example of multi-modal transfer learning between DNA and protein language models for the task of predicting transcript isoform expression. These recent successes suggest that fine-tuned gLMs may make meaningful progress on diverse genome interpretation tasks.
Two recent studies evaluated several gLMs in prediction tasks in the human genome and found that they generally did not outperform non-gLM baselines [50, 51]. These results were based on frozen embeddings; evaluating full fine-tuning would provide additional insights. While gLMs are already well suited to demonstrate the value of transfer learning in less-studied organisms, further innovation may be required for them to offer significant value in human genetics, where high-quality labeled data and carefully crafted models already exist. An important question is how far the scaling hypothesis holds for gLMs, i.e., how much increasing unlabeled data and computation will keep improving model performance. A recent pLM study found that scaling improved only protein structure prediction but not most other tasks such as function or property prediction [52], so gLM tasks should also be held to the same scrutiny.
@@
LookingGlass [58] Bacteria + archaea CLM Recurrent Neural Network nucleotide-level Metagenomic sequences from diverse environments rather than assembled genomes are used for training.
@@
SpliceBERT [64] Human + 71 vertebrate genomes MLM Transformer nucleotide-level Only RNA Transcripts are used in training.
SpeciesLM Fungi [65] 1500 fungal genomes MLM Transformer overlapping $k$-mer Only $5^{\prime}$ and $3^{\prime}$ UTR regions are used in training: the $5^{\prime}$ species LM and $3^{\prime}$ species LM.
@@
UTR-LM [69] Human + 4 vertebrate genomes MLM Transformer nucleotide-level Only $5^{\prime}$ UTR regions are used in training. Trained also to predict mRNA minimum free energy and secondary structures calculated by ViennaRNA [70].
@@
AgroNT [14] 48 plant genomes focusing on edible plant species MLM Transformer non-overlapping $k$-mer
@@
regLM [30] Human + yeast CLM SSM nucleotide-level Human enhancer and yeast promoter sequences are used to fine-tune/pretrain separate HyenaDNA [35] models
@@
ChatNT [74] Genomic sequences + English instructions CLM Transformer overlapping $k$-mer Combines the pretrained gLM NT [16] and the English LM Vicuna [75]. Trained to perform all supervised genomics prediction tasks as text-to-text tasks.
LucaOne [76] Genomic and protein sequences from 169,861 species MLM Transformer nucleotide- and amino acid-level Mixed pretraining with DNA, RNA, and protein sequences. Trained also to predict 8 types of selected annotations.
@@
CD-GPT [77] Genomic and protein sequences of 14 organisms CLM Transformer BPE Mixed pretraining with DNA, RNA, and protein sequences, followed by targeted DNA-Protein and mRNA-Protein paired pretraining.
SpeciesLM Metazoa [19] 494 metazoan genomes MLM Transformer overlapping $k$-mer Only trained on 2 kb upstream of start codons
gLM2 [78] Metagenomes and genomes from IMG [79] and MGnify [80]. MLM Transformer BPE for nucleotides, amino acid-level for proteins Pretraining with a mixed-modality dataset, comprising interleaved protein-coding (amino acid) and intergenic (nucleotide) sequences.
Table 1: A summary of existing gLMs. An overview of various gLMs is provided, highlighting their pretraining datasets, tasks, architectures, tokenization methods, and unique features. The models are listed in the order of their public release dates. Abbreviations used include SSM for State Space Model, CNN for Convolutional Neural Network, BPE for Byte-Pair Encoding, CLM for Causal Language Modeling, and MLM for Masked Language Modeling.
@@
The performance of a machine learning model is significantly influenced by both its architecture and its training data. Various model architectures such as convolutional neural networks (CNNs), Transformers, and state-space models have been successfully adapted to a wide range of domains, including natural language, images, audio, proteins, and genomics. However, selecting suitable data for pretraining requires a deep understanding of the specific domain, especially in genomics where there is no universally accepted, curated dataset comparable to those in NLP (e.g., the Pile [81, 82]) or protein biology (e.g., UniProt [4]).
@@
Another key question is how to ensure that the amount of data is enough. It is likely that a single genome might not be enough to train a large model, especially if non-functional regions are downsampled or downweighted. One approach is to add sequence variants from the same species [16]. However, in many species, including humans, there is relatively little variation between individuals. A more common approach is to train across multiple species (Table 1), as typically done for pLMs. As species become more distant, regulatory logic diverges faster than proteins. One proposed approach is to explicitly add a species identifier as an extra input to the model [65]. Notwithstanding, it is plausible that a large enough model, with enough genomic context, might be able to naturally model distant genomes, similarly to how LLMs handle multilingual datasets.
As mentioned earlier, in prokaryotes, there exist models (MegaDNA and EVO) that take an entire genome as context [36, 31]. This is currently infeasible for eukaryotes, and therefore leads to the question of how to partition the genome into context windows to be separately modeled. Many interactions are restricted to nearby positions, such as transcription factor binding site motifs, motivating the development of models with a relatively small context ($<$ 6 kb) (Table 1). However, there are obvious long-range interactions, such as between exons of the same gene or between enhancers and promoters (up to 1 Mb) [85]. Such long context lengths introduce computational and statistical challenges, and efforts have been made to overcome them [35, 36, 31, 24]. Regardless of the chosen context length, it is still not easy to partition the genome into independent units (similarly to how proteomes are separated by protein). For instance, the enhancer of a gene can be located inside the intron of another gene [85], and multiple genes can be controlled by the same enhancer [86]. Avoiding potential data leakage due to orthology and paralogy is quite challenging, especially when training across species.
@@
Despite their strengths, Transformer models face several challenges unique to genomic modeling. One significant issue is that Transformers have weak or no inductive biases regarding the locality of interactions [91, 53], making them less data-efficient at modeling local motifs such as TFBS. This motivated the development of CNN-Transformer hybrids such as LOGO [59], following supervised models such as Enformer [22].
@@
To overcome the quadratic scaling issues of self-attention, various state-space models (SSMs) [94, 95, 96] have been proposed for gLMs as efficient alternatives to Transformers, offering nearly linear scaling with sequence length. HyenaDNA [35], based on the Hyena Hierarchy [95], can support input contexts as long as 1 million nucleotides. EVO [31], a hybrid model combining Hyena and Transformer architectures, is pretrained with 8 kb sequences and later fine-tuned with 131 kb sequences during the context extension stage. Caduceus [24], built on the Mamba-based SSM [96], is trained on 131 kb sequences while incorporating reverse-complement equivariance.
@@
To reduce input sequence length and model longer context, both $k$-mer and byte-pair encoding [101] (BPE) tokenizations create artificially defined nucleotide vocabularies larger than the natural nucleotide vocabularies of {A, C, G, T}. On the other hand, single-nucleotide tokenization simplifies model interpretation and attribution, and enhances the model’s ability to handle genomic variations more effectively.
Several modifications to the training objective have been explored to provide additional signal and boost performance. For instance, GPN-MSA [17] enhances MLM training on the human reference genome with a whole-genome MSA [102, 103] of vertebrate species, leveraging conservation across related species for additional context. A limitation is that whole-genome MSAs have only been generated for certain species, and might require further development to be effective in plants [104]. Additionally, cis regulatory elements tend to diverge fast in sequence space even if they have conserved activity, which limits the orthology information that can be extracted via alignment [105]. Species LM [65] directly integrates species information by assigning a dedicated token for each yeast species and appending the species token to the input sequence during training and inference. Pretraining on nucleotide sequences has been expanded to enable cross-talk with additional modalities such as epigenetics [55, 56], RNA [77, 76], proteins [77, 78, 76], and natural language [74].